# Another look at inference after prediction

### v2

@@ 3.2 Asymptotic analysis of the CC estimator
To further analyze the efficiency of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ we draw on a landmark paper that introduced a class of augmented inverse probability weighted (AIPW) estimators which includes all regular asymptotic linear estimators (Robins et al., 1994). In the current setting, where the goal is to estimate linear regression coefficients $\bm{\beta}^{*}$ with $Y$ partially observed, but $\widehat{Y}$ fully observed, these estimators have influence functions of the form
@@
By Theorem 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is an AIPW estimator and can be seen as an approximation to the optimal estimator with $\bm{\omega}(\widehat{Y},{\bf{X}})=\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$. The PPI estimator is therefore not the most efficient estimator in its class and, from Corollary 1.1, can even be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\boldsymbol{\Sigma}_{Y,\widehat{Y}}$. A PB inference strategy should enhance statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever possible and, at the very least, should not be negatively impacted by incorporating unlabeled data into estimation. This raises concerns about the practical utility of the PPI estimator.
To improve upon $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, one approach is to assume a parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ that can be estimated from $\mathscr{L}$ for analysis. However, it is often difficult to specify an appropriate parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$, particularly when the relationship between $Y$ and $\widehat{Y}$ is unknown. Moreover, when the model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ is misspecified, an efficiency gain is not guaranteed. To overcome these difficulties, Chen and Chen (2000) proposed a particular choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ that is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that does not require modeling of $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$. Their approach is motivated by the double-sampling design which, apart from differences in terminology and notation (i.e., they regard $\widehat{Y}$ as a mismeasured version of $Y$), is analogous to the partially labeled data setting considered here. The CC estimator approximates $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ with
@@
obtained with $\mathscr{L}$. Therefore, a simple way to ensure that an augmented estimator like the PPI estimator has asymptotic variance at least as low as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is to add a weight function in front of the augmentation term. In fact, the framework of Robins et al. (1994) implies that all regular asymptotically linear estimators for $\bm{\beta}^{*}$ augment $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with an augmentation term that incorporates information from $\widehat{Y}$. We present the influence function expansion for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ in Theorem 2 to further study its asymptotic variance and compare it with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$. All of the asymptotic results provided for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ are detailed in Appendix C.
Theorem 2 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$)
Following Chen and Chen (2000),
@@
Corollary 2.1 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$)
@@
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq{\bf{0}}.$
@@
Corollary 2.2 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is the most efficient estimator in its class)
Let the mean squared error (MSE) of an estimator $\hat{\bm{\beta}}$ for $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$. Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum (2015), page 40), ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
From Corollary 2.1, we note that in contrast to $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$, the covariance term, $\boldsymbol{\Sigma}_{Y,\widehat{Y}}$, in $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ is normalized by the residual variance of $\widehat{Y}$, $\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$. This normalization effectively scales the covariance according to the precision of $\widehat{Y}$ and enables $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ to always be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Coming back to the example from the previous section, we can gain more intuition as to why $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ possesses this property while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ does not.
@@ 4 Related literature
Existing PB inference literature has established connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error (Angelopoulos et al., 2023a, b; Miao et al., 2023; Zrnic and Candès, 2024a, b). Here we draw connections between PB inference methods and methods from the economics and statistics literature dating back to the 1960s. These connections reveal that aspects of the PB inference problem have been explored in various areas, albeit with differences in motivation, terminology, and notation. Given the close relationship between the PPI and CC methods, future research in PB inference could benefit from a more comprehensive review of these and other related approaches than what is provided here.
@@ 4.1 Seemingly unrelated regression & its use in statistical genetics
Seemingly unrelated regression (SUR) was introduced by economist Arnold Zellner in 1962 for simultaneous estimation of multiple related regression models (Zellner, 1962; Swamy and Mehta, 1975; Schmidt, 1977). The regression models have different outcome variables, but are assumed to have correlated errors. SUR leverages the correlation of the error terms to produce more efficient estimators compared to estimating each regression independently. In the current setting, the SUR problem can be stated formally as
@@
where $\mathbb{E}(\epsilon\mid{\bf{X}})=0$ and $\mathbb{E}(\epsilon^{\prime}\mid{\bf{X}})=0.$ An estimator for $\bm{\beta}^{*}$ that is based on augmenting $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ was proposed by Conniffe (1985) and is given by:
@@
Note that $\widehat{W}^{\text{\tiny{SUR}}}$ is an estimator of the scalar $W^{\text{\tiny{SUR}}}=\frac{\mathbb{E}[\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})]}{\mathbb{E}[\text{Var}(\epsilon^{\prime}\mid{\bf{X}})]}$. Similar to the PPI and CC methods, the contribution of the augmentation term to estimation increases with the predictiveness of $\widehat{Y}$ for $Y$. In the extreme case where $\widehat{Y}$ is a perfect predictor for $Y$ such that $\widehat{Y}=Y$, $\widehat{W}^{\text{\tiny{SUR}}}=1$ and $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ is simply $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$. In the homoskedastic case when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant and do not depend on ${\bf{X}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (Chen and Chen, 2000). Moreover, when the conditional distribution of $(\epsilon,\epsilon^{\prime})$ given ${\bf{X}}$ is bivariate normal with mean $\boldsymbol{0}$ and has constant variance-covariance, $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (and thus $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ as well) is the maximum likelihood estimator of $\bm{\beta}^{*}$ and fully efficient under the assumed parametric model (Conniffe, 1985). In other scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves a lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$.
The SUR estimator was recently applied for a PB inference problem inspired by genome-wide association studies conducted with population biobanks. These studies often rely on pre-trained models to impute outcomes of interest (i.e., phenotypes or traits) as they are partially missing due to the time and expense of ascertainment (McCaw et al., 2024). The proposed method, “SynSurr”, assumes the true and predicted outcomes follow a bivariate normal distribution. The SynSurr estimator is equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ and asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant. Building on SynSurr, Miao et al. (2024) proposed the “POP-GWAS” estimator for a variety of regression models utilizing the augmentation technique. The POP-GWAS estimator is
@@ 4.2 Augmented estimation for quantifying a treatment effect
Outside of statistical genetics, the augmented estimation approach used by the PPI and CC methods has been widely used for treatment effect estimation in the statistics and biostatistics literature (e.g., Leon et al. (2003); Lu and Tsiatis (2008); Tsiatis et al. (2008); Gilbert et al. (2009); Zhang and Gilbert (2010); Tian et al. (2012); Yang and Ding (2020)). Many of these approaches build on methods from the survey sampling literature (e.g., Cassel et al. (1976); Särndal et al. (2003)).
For example, Tsiatis et al. (2008) consider covariate adjusted analyses for comparing two treatments in randomized clinical trials. Letting $X$ be an indicator for randomization to the treated or control arms, the focus of inference is on $\beta$ in the model
@@
in order to estimate the treatment effect $\Delta=\mathbb{E}(Y\mid X=1)-\mathbb{E}(Y\mid X=0)$. A standard estimator of $\Delta$ is $\widehat{\Delta}=\overline{Y}^{1}-\overline{Y}^{0}$, where $\overline{Y}^{1}=n_{1}^{-1}\sum_{i=1}^{n}Y_{i}X_{i}$ and $\overline{Y}^{0}=n_{0}^{-1}\sum_{i=1}^{n}Y_{i}(1-X_{i})$ are the empirical means of $Y$ in the treated and control groups of sizes $n_{1}$ and $n_{0}$, respectively. To utilize information in baseline covariates, ${\bf{Z}}$, Tsiatis et al. (2008) propose an augmented estimator of the treatment effect given by

### v3

@@ 3.2 Asymptotic analysis of the CC estimator
To further analyze the efficiency of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ we draw on the landmark paper that introduced a class of augmented inverse probability weighted (AIPW) estimators which includes all regular asymptotic linear estimators [Robins et al., 1994]. In the current setting, where the goal is to estimate linear regression coefficients $\bm{\beta}^{*}$ with $Y$ partially observed, but $\widehat{Y}$ fully observed, these estimators have influence functions of the form
@@
By Theorem 1, $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is an AIPW estimator and can be seen as an approximation to the optimal estimator with $\bm{\omega}(\widehat{Y},{\bf{X}})=\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})$. The PPI estimator is therefore not the most efficient estimator in its class and, from Corollary 1.1, can even be less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ when $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\bm{\Sigma}_{Y,\widehat{Y}}$. A PB inference strategy should enhance statistical efficiency relative to $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ whenever possible and, at the very least, should not be negatively impacted by incorporating unlabeled data into estimation. This raises concerns about the practical utility of the PPI estimator.
To improve upon $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$, one approach is to assume a parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ that can be estimated from $\mathscr{L}$ for analysis. However, it is often difficult to specify an appropriate parametric model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$, particularly when the relationship between $Y$ and $\widehat{Y}$ is unknown. Moreover, when the model for $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ is misspecified, an efficiency gain is not guaranteed. To overcome these difficulties, Chen and Chen [2000] proposed a particular choice of $\bm{\omega}(\widehat{Y},{\bf{X}})$ that is guaranteed to be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ and that does not require modeling of $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$. Their approach is motivated by the double-sampling design which, apart from differences in terminology (i.e., they regard $\widehat{Y}$ as a mismeasured version of $Y$), is analogous to the partially labeled data setting considered here. The CC estimator approximates $\bm{\omega}^{\text{\tiny{opt}}}(\widehat{Y},{\bf{X}})$ with
@@
obtained with $\mathscr{L}$. Therefore, a simple way to ensure that an estimator like the PPI estimator has asymptotic variance at least as low as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is to add a weight function in front of the augmentation term. In fact, the framework of Robins et al. [1994] implies that all regular asymptotically linear estimators for $\bm{\beta}^{*}$ augment $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ with an augmentation term that incorporates information from $\widehat{Y}$. We present the influence function expansion for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ in Theorem 2 to further study its asymptotic variance and compare it with $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$. All of the asymptotic results provided for $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ are detailed in Appendix C.
Theorem 2 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$).
Following Chen and Chen [2000],
@@
Corollary 2.1 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
@@
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq{\bf{0}}.$
@@
Corollary 2.2 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is the most efficient estimator in its class).
Let the mean squared error (MSE) of an estimator $\hat{\bm{\beta}}$ for $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$. Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum [2015], page 40), ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
From Corollary 2.1, we note that in contrast to $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}})$, the covariance term, $\bm{\Sigma}_{Y,\widehat{Y}}$, in $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})$ is normalized by the residual variance of $\widehat{Y}$, $\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$. This normalization effectively scales the covariance according to the precision of $\widehat{Y}$ and enables $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ to always be at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. Coming back to the example from the previous section, we can gain more intuition as to why $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ possesses this property while $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ does not.
@@ 4 Related literature
Existing PB inference literature has established connections with semi-supervised inference, semi-parametric inference, missing data, and measurement error [Angelopoulos et al., 2023a, b, Miao et al., 2023, Zrnic and Candès, 2024a, b, Angelopoulos et al., 2024]. Here we draw connections between PB inference methods and modern and classical methods from the economics and statistics literature dating back to the 1960s. These connections reveal that aspects of the PB inference problem have been explored in various areas, albeit with differences in motivation, terminology, and notation. Given the close relationship between the PPI and CC methods, future research in PB inference could benefit from a more comprehensive review of these and other related approaches than what is provided here.
@@ 4.1 Seemingly unrelated regression & its use in statistical genetics
Seemingly unrelated regression (SUR) was introduced by economist Arnold Zellner in 1962 for simultaneous estimation of multiple related regression models [Zellner, 1962, Swamy and Mehta, 1975, Schmidt, 1977]. The regression models have different outcome variables, but are assumed to have correlated errors. SUR leverages the correlation of the error terms to produce more efficient estimators compared to estimating each regression independently. In the current setting, the SUR problem can be stated formally as
@@
where $\mathbb{E}(\epsilon\mid{\bf{X}})=0$ and $\mathbb{E}(\epsilon^{\prime}\mid{\bf{X}})=0.$ An estimator for $\bm{\beta}^{*}$ that is based on augmenting $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ was proposed by Conniffe [1985] and is given by:
@@
Note that $\widehat{W}^{\text{\tiny{SUR}}}$ is an estimator of the scalar $W^{\text{\tiny{SUR}}}=\frac{\mathbb{E}[\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})]}{\mathbb{E}[\text{Var}(\epsilon^{\prime}\mid{\bf{X}})]}$. Similar to the PPI and CC methods, the contribution of the augmentation term to estimation increases with the predictiveness of $\widehat{Y}$ for $Y$. In the extreme case where $\widehat{Y}$ is a perfect predictor for $Y$ such that $\widehat{Y}=Y$, $\widehat{W}^{\text{\tiny{SUR}}}=1$ and $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ is simply $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$. In the homoskedastic case when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant and do not depend on ${\bf{X}}$, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ [Chen and Chen, 2000]. Moreover, when the conditional distribution of $(\epsilon,\epsilon^{\prime})$ given ${\bf{X}}$ is bivariate normal with mean $\bm{0}$ and has constant variance-covariance, $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ (and thus $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ as well) is the maximum likelihood estimator of $\bm{\beta}^{*}$ and fully efficient under the assumed parametric model [Conniffe, 1985]. In other scenarios, $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ achieves a lower asymptotic variance than $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$.
The SUR estimator was recently applied for a PB inference problem inspired by genome-wide association studies conducted with population biobanks. These studies often rely on pre-trained models to impute outcomes of interest (i.e., phenotypes or traits) as they are partially missing due to the time and expense of ascertainment [McCaw et al., 2024]. The proposed method, “SynSurr”, assumes the true and predicted outcomes follow a bivariate normal distribution. The SynSurr estimator is equivalent to $\hat{\bm{\beta}}^{\text{\tiny{SUR}}}$ and asymptotically equivalent to $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ when $\text{Var}(\epsilon\mid{\bf{X}})$ and $\text{Cov}(\epsilon,\epsilon^{\prime}\mid{\bf{X}})$ are constant. Building on SynSurr, Miao et al. [2024] proposed the “POP-GWAS” estimator for a variety of regression models utilizing the augmentation technique. The POP-GWAS estimator is
@@ 4.2 Augmented estimation for quantifying a treatment effect
Outside of statistical genetics, the augmented estimation approach used by the PPI and CC methods has been widely used for treatment effect estimation in the statistics and biostatistics literature (e.g., Leon et al. [2003], Lu and Tsiatis [2008], Tsiatis et al. [2008], Gilbert et al. [2009], Zhang and Gilbert [2010], Tian et al. [2012], Yang and Ding [2020]). Many of these approaches build on methods from the survey sampling literature (e.g., Cassel et al. [1976], Särndal et al. [2003]).
For example, Tsiatis et al. [2008] consider covariate adjusted analyses for comparing two treatments in randomized clinical trials. Letting $X$ be an indicator for randomization to the treated or control arms, the focus of inference is on $\beta$ in the model
@@
in order to estimate the treatment effect $\Delta=\mathbb{E}(Y\mid X=1)-\mathbb{E}(Y\mid X=0)$. A standard estimator of $\Delta$ is $\widehat{\Delta}=\overline{Y}^{1}-\overline{Y}^{0}$, where $\overline{Y}^{1}=n_{1}^{-1}\sum_{i=1}^{n}Y_{i}X_{i}$ and $\overline{Y}^{0}=n_{0}^{-1}\sum_{i=1}^{n}Y_{i}(1-X_{i})$ are the empirical means of $Y$ in the treated and control groups of sizes $n_{1}$ and $n_{0}$, respectively. To utilize information in baseline covariates, ${\bf{Z}}$, Tsiatis et al. [2008] propose an augmented estimator of the treatment effect given by